# Novel Strontium Carbides Under Compression

### v2

@@ III Results
Sr2C5 crystallizes in the low-symmetry space group $P\bar{1}$. C atoms in this structure feature extended ribbons consisting of C-pentagons and C-hexagons, as shown in Fig. 4(g, h).The C-C bond lengths range from 1.39 Å to 1.44 Å, indicating an intermediate bond order between single and double C-C bonds. In the Sr2C5 structure, one Sr atom has +1.248e Bader charge and is 13-fold coordinated, while another Sr atom has +1.202e Bader charge and is 12-fold coordinated (see Fig. 7(a)). In the fused C-hexagon and C-pentagon, carbon atoms are negatively charged. Therefore, it is evident that there is charge transfer from Sr to C atoms. Sr atom with 13-fold coordination is surrounded by the C-hexagon and a heptagon, while the Sr atom with 12-fold coordination is surrounded by the C-pentagon and a heptagon. In the C-hexagons, the four inner C atoms have three C-C covalent bonds each, while the two outer C atoms have two C-C covalent bonds and one lone pair of electrons each. The C-C bonds in the C-hexagon form two different angles $\alpha$ and $\beta$, where $\alpha$=112 and $\beta$= 124. In C-pentagons, the inner four C atoms have three C-C covalent bonds, while the outer C atom has two C-C covalent bonds and a lone pair of electrons. The lone pairs of electrons can be seen in the 2-D electron localization function (see Fig. 7 (b)).
@@
Finally, SrC3 crystallizes in the $Pnma$ space group. In this structure C atoms form infinite ribbons consisting of pentagonal carbon rings (see SM Fig. S4 (a)). Bader analysis indicates a charge transfer from strontium to carbon atoms, with strontium atoms acquiring a positive Bader charge of +1.343 (see SM Tab. 1). The carbon atoms, in contrast, have negative charges, with those not involved in ring fusion showing higher electron concentration. This is also evident in the 2D ELF plot (see SM Fig. S4 (b)).
@@
In light of the appreciable diversity of the observed carbon motifs (see Figure 4), it is instructive to check if at least some of these configurations can be explained using conventional chemical wisdom. For that purpose, we investigated the applicability of the Zintl-Klemm (ZK) concept, originally conceived to rationalize bonding patterns of polyanions, to the newly identified strontium carbides. To assess the validity of the ZK rules, the average charge of a C-anion will be determined by assigning a formal $+2$ charge to the Sr atoms (Sr2+) and imposing charge neutrality of the formal SrxCy unit. For the example of Sr2C this would lead to a formal charge of the carbon anion of $-4$. By subsequently equating the structural motifs formed by that C-anion to the isoelectronic, neutral species, in the case of C4- that would be a neon atom, the structure of the polyanion is inferred. Since neon is a noble gas, and Ne-Ne bonds cannot be formed, we expect no C-C bonds in Sr2C, which we indeed observe. We performed this analysis for all herein discussed stoichiometries, which we have summarized in Table 1. We find that for some of the strontium carbide compounds, namely Sr2C, Sr3C2 and SrC, the C-anion structure is consistent with the motifs predicted from the ZK rule. This finding is also in agreement with a previous investigation of high-pressure Ca-C phases Li et al. (2015). In the other cases, however, we find that the ZK concept cannot be directly applied, since for some stoichiometries including Sr2C3, SrC3 and Sr2C5, the average formal charge on a C-atom would be fractional. Double and triple bonds also lead to exceptions from the ZK rule. A particularly noteworthy example is the case of SrC2, in which the carbon species would be isoelectronic to a nitrogen atom. As a consequence one would expect each carbon atom to form three covalent bonds. As illustrated in Figure 4, we find that the carbon polyanion in $Immm$-SrC2 consists of planar, napthalene-like arrangements. The planarity and the similarity of all C-C bond lengths in that polyanion suggests that that configuration is aromatic. This implies that most of the carbon atoms form one single and one double bond, so that the bonding of these atoms in $Immm$-SrC2 is consistent with the ZK prediction. The two carbon atoms shared between the two rings, however, must then consequently have two single bonds and one double bond, which is not compatible with the isoelectronic N-atom. The same argument applies to the planar structure of the anion in the $P\overline{1}$ phase of SrC2, so that both SrC2 phases are in contradiction with the ZK rule.
@@ V Acknowledgements
N.R. and A.S. acknowledge funding from the Russian Science Foundation (Project No. 23-13-00332). A.R.O. and P.G. acknowledge funding from the Russian Science Foundation (Project No. 19-72-30043). E.M. thanks the NOMAD Laboratory at the FHI of the Max Planck Society for support of the project.

### v3

@@ III Results
Sr2C5 crystallizes in the low-symmetry space group $P\bar{1}$. C atoms in this structure feature extended ribbons consisting of C-pentagons and C-hexagons, as shown in Fig. 4(g, h).The C-C bond lengths range from 1.39 Å to 1.44 Å, indicating an intermediate bond order between single and double C-C bonds. In the Sr2C5 structure, one Sr atom has +1.248 Bader charge and is 13-fold coordinated, while another Sr atom has +1.202 Bader charge and is 12-fold coordinated (see Fig. 7(a)). In the fused C-hexagon and C-pentagon, carbon atoms are negatively charged. Therefore, it is evident that there is charge transfer from Sr to C atoms. Sr atom with 13-fold coordination is surrounded by the C-hexagon and a heptagon, while the Sr atom with 12-fold coordination is surrounded by the C-pentagon and a heptagon. In the C-hexagons, the four inner C atoms have three C-C covalent bonds each, while the two outer C atoms have two C-C covalent bonds and one lone pair of electrons each. The C-C bonds in the C-hexagon form two different angles $\alpha$ and $\beta$, where $\alpha$=112 and $\beta$= 124. In C-pentagons, the inner four C atoms have three C-C covalent bonds, while the outer C atom has two C-C covalent bonds and a lone pair of electrons. The lone pairs of electrons can be seen in the 2-D electron localization function (see Fig. 7 (b)).
@@
Finally, SrC3 crystallizes in the $Pnma$ space group. In this structure, C atoms form infinite C-ribbons consisting of pentagonal carbon rings (see SM Fig. S4 (a)). Bader analysis indicates a charge transfer from strontium to carbon atoms, with strontium atoms acquiring a positive Bader charge of +1.343. The carbon atoms, in contrast, have negative charges, with those not involved in ring fusion showing higher electron concentration. This is also evident in the 2D ELF plot (see SM Fig. S4 (b)).
@@
In light of the appreciable diversity of the observed carbon motifs as demonstrated in Fig. 4, it is instructive to check whether at least some of these configurations can be explained using conventional chemical wisdom. For that purpose, we investigated the applicability of the Zintl-Klemm (ZK) concept, originally conceived to rationalize bonding patterns of polyanions, to the newly identified strontium carbides. To assess the validity of the ZK rules, the average charge of a C-anion will be determined by assigning the formal $+2$ charge to the Sr atoms (Sr2+) and imposing charge neutrality of the formal SrxCy unit. For the example of Sr2C this would lead to a formal charge of the carbon anion of $-4$. By subsequently equating the structural motifs formed by that C-anion to the isoelectronic, neutral species, in the case of C4- that would be a neon atom, the structure of the polyanion is inferred. Since neon is a noble gas, and Ne-Ne bonds cannot be formed, we expect no C-C bonds in Sr2C, which we indeed observe. We performed this analysis for all herein discussed stoichiometries, which we have summarized in Tab. 1. We find that for some of the strontium carbide compounds, namely Sr2C, Sr3C2 and SrC, the C-anion structure is consistent with the motifs predicted from the ZK rule. This finding is also in agreement with a previous investigation of high-pressure Ca-C phases Li et al. (2015). In the other cases, however, we find that for some stoichiometries such as SrC3 and Sr2C5 the ZK concept cannot be directly applied. Double and triple bonds also lead to exceptions from the ZK rule. A particularly noteworthy example is the case of SrC2, in which the carbon species would be isoelectronic to a nitrogen atom. As a consequence, one would expect each carbon atom to form three covalent bonds. As illustrated in Fig. 4, we find that the carbon polyanion in $Immm$-SrC2 consists of planar, napthalene-like arrangements. The planarity and the similarity of all C-C bond lengths in that polyanion suggests that the configuration is aromatic. This implies that most of the carbon atoms form one single and one double bond, so that the bonding of these atoms in $Immm$-SrC2 is consistent with the ZK prediction. The two carbon atoms shared between the two rings, however, must then consequently have two single bonds and one double bond, which is not compatible with the isoelectronic N-atom. The same argument applies to the planar structure of the anion in the $P\overline{1}$ phase of SrC2, so that both SrC2 phases are in contradiction with the ZK rule.
@@ V Acknowledgements
N.R. and A.S. acknowledge funding from the Russian Science Foundation (Project No. 23-13-00332). A.R.O. and P.G. acknowledge funding from the Russian Science Foundation (Project No. 19-72-30043). E.M. thanks the NOMAD Laboratory at the FHI of the Max Planck Society.